# SMEFT Matching to $Z^{\prime}$ Models at Dimension-8

### v2

@@ II Models with a heavy $Z^{\prime}$
$l_{L}^{k}$ $-\frac{1}{2}$ $-1$ $\delta_{ik}-\delta_{jk}$ $\frac{1}{2\sqrt{6}}$ $\frac{3}{2\sqrt{10}}$ $\frac{1}{2\sqrt{15}}$ $-\frac{1}{2}$ $\frac{1}{\sqrt{10}}$ $\frac{1}{\sqrt{15}}$
$\ell^{k}_{R}$ $-1$ $-1$ $\delta_{ik}-\delta_{jk}$ $\frac{1}{2\sqrt{6}}$ $\frac{-1}{2\sqrt{10}}$ $\frac{-1}{\sqrt{15}}$ 00 $\frac{1}{2\sqrt{10}}$ $\frac{-1}{\sqrt{15}}$
@@ III SMEFT and Matching
$\mathcal{L}_{Z^{\prime}}=-\frac{1}{4}Z^{\prime}_{\mu\nu}Z^{\prime\mu\nu}+\frac{1}{2}M_{Z^{\prime}}^{2}Z^{\prime}_{\mu}Z^{\prime\mu}-\frac{\epsilon}{2}B_{\mu\nu}Z^{\prime\mu\nu}+\left(g_{H,2}\right)^{2}Z^{\prime}_{\mu}Z^{\prime\mu}|H^{\dagger}H|-Z^{\prime}_{\mu}\mathcal{J}^{\mu},$ (4)
@@
where $H^{\dagger}\overleftrightarrow{D}^{\mu}H=H^{\dagger}(D^{\mu}H)-(D^{\mu}H^{\dagger})H$ with $D_{\mu}=\partial_{\mu}+ig^{\prime}Y_{H}B_{\mu}+igW^{a}_{\mu}T^{a}$ the usual covariant derivative. The couplings of Eqs. 4 and (5) can be determined for each of the models described in the previous section in terms of the $U(1)^{\prime}$ gauge coupling $g_{D}$ and the particle charges of Table 1,
@@
$\begin{split}\delta\mathcal{L}=&-\frac{1}{2M_{Z^{\prime}}^{2}}\mathcal{J}^{\mu}\mathcal{J}_{\mu}-\frac{\epsilon}{M_{Z^{\prime}}^{2}}\left(\partial_{\nu}B^{\mu\nu}\right)\mathcal{J}_{\mu}-\frac{\epsilon^{2}}{2M_{Z^{\prime}}^{2}}\left(\partial_{\nu}B^{\mu\nu}\right)^{2}\\ &+\frac{\epsilon^{2}}{2M_{Z^{\prime}}^{4}}\left(\partial_{\nu}B^{\mu\nu}\right)\partial^{2}\left(\partial^{\alpha}B_{\mu\alpha}\right)+\frac{\epsilon}{M_{Z^{\prime}}^{4}}\left(\partial_{\nu}B^{\mu\nu}\right)\partial^{2}\mathcal{J}_{\mu}+\frac{1}{2M_{Z^{\prime}}^{4}}\mathcal{J}^{\mu}\partial^{2}\mathcal{J}_{\mu}\\ &+\frac{\epsilon^{2}g_{H,2}^{2}}{M_{Z^{\prime}}^{4}}H^{\dagger}H\left(\partial_{\nu}B^{\mu\nu}\right)^{2}+\frac{2\epsilon g_{H,2}^{2}}{M_{Z^{\prime}}^{4}}H^{\dagger}H\left(\partial_{\nu}B^{\mu\nu}\right)\mathcal{J}_{\mu}+\frac{g_{H,2}^{2}}{M_{Z^{\prime}}^{4}}H^{\dagger}H\mathcal{J}^{\mu}\mathcal{J}_{\mu}\,,\end{split}$ (9)
@@
$\begin{split}B_{\mu}&\rightarrow B_{\mu}-\frac{\epsilon^{2}}{2M_{Z^{\prime}}^{2}}\left[\left(\partial^{\nu}B_{\mu\nu}\right)+j_{\mu}\right]-\frac{\epsilon}{M_{Z^{\prime}}^{2}}\mathcal{J}_{\mu}\,,\\ \partial_{\nu}B^{\mu\nu}&\rightarrow\partial_{\nu}B^{\mu\nu}+\frac{\epsilon^{2}}{2M_{Z^{\prime}}^{2}}\partial^{2}\left(\partial_{\nu}B^{\mu\nu}\right)+\frac{\epsilon^{2}}{2M_{Z^{\prime}}^{2}}\partial^{2}j^{\mu}+\frac{\epsilon}{M_{Z^{\prime}}^{2}}\partial^{2}\mathcal{J}^{\mu},\end{split}$ (10)
where $j_{\mu}=\frac{ig^{\prime}}{2}\left(H^{\dagger}\overleftrightarrow{D^{\mu}}H\right)+g^{\prime}\sum_{f}Y_{f}\bar{f}\gamma^{\mu}f$ is the SM hypercharge current with $f\in(Q_{L},L_{L},u_{R},d_{R},e_{R})$, $g^{\prime}$ is the SM hypercharge gauge coupling and we omit generation indices for simplicity. Including the newly generated dimension- 8 terms, this gives us
$\begin{split}\delta\mathcal{L}=&-\frac{1}{2M_{Z^{\prime}}^{2}}\mathcal{J}^{\mu}\mathcal{J}_{\mu}-\frac{\epsilon^{2}}{2M_{Z^{\prime}}^{2}}j^{\mu}j_{\mu}-\frac{\epsilon}{M_{Z^{\prime}}^{2}}j^{\mu}\mathcal{J}_{\mu}\\ &+\frac{1}{M_{Z^{\prime}}^{4}}\bigg{\{}\left(\frac{\epsilon^{2}}{2}-\frac{3\epsilon^{4}}{8}\right)\left(\partial_{\nu}B^{\mu\nu}\right)\partial^{2}\left(\partial^{\alpha}B_{\mu\alpha}\right)+\left(\epsilon-\epsilon^{3}\right)\left(\partial_{\nu}B^{\mu\nu}\right)\partial^{2}\mathcal{J}_{\mu}\\ &-\frac{\epsilon^{4}}{4}\left(\partial_{\nu}B^{\mu\nu}\right)\partial^{2}j_{\mu}+\left(\frac{1}{2}-\frac{\epsilon^{2}}{2}\right)\mathcal{J}_{\mu}\partial^{2}\mathcal{J}^{\mu}+\frac{\epsilon^{4}}{8}j^{\mu}\partial^{2}j_{\mu}+\frac{g^{\prime 2}\epsilon^{3}}{4}\left(H^{\dagger}H\right)\mathcal{J}^{\mu}j_{\mu}\\ &+\left(\epsilon^{2}g_{H,2}^{2}+\frac{g^{\prime 2}\epsilon^{4}}{16}\right)\left(H^{\dagger}H\right)\left(\partial_{\nu}B^{\mu\nu}\right)^{2}+\left(2\epsilon g_{H,2}^{2}+\frac{g^{\prime 2}\epsilon^{3}}{4}\right)\left(H^{\dagger}H\right)\left(\partial_{\nu}B^{\mu\nu}\right)\mathcal{J}_{\mu}\\ &+\left(g_{H,2}^{2}+\frac{g^{\prime 2}\epsilon^{2}}{4}\right)\left(H^{\dagger}H\right)\mathcal{J}^{\mu}\mathcal{J}_{\mu}+\frac{g^{\prime 2}\epsilon^{4}}{16}\left(H^{\dagger}H\right)j_{\mu}j^{\mu}+\frac{g^{\prime 2}\epsilon^{4}}{8}\left(H^{\dagger}H\right)\left(\partial_{\nu}B^{\mu\nu}\right)j_{\mu}\bigg{\}}\,.\end{split}$ (11)
@@
$\begin{split}\delta\mathcal{L}=&-\frac{1}{2M_{Z^{\prime}}^{2}}\left(\mathcal{J}_{\mu}+\epsilon j_{\mu}\right)^{2}\\ &-\frac{1}{2M_{Z^{\prime}}^{4}}\left(1-\epsilon^{2}\right)\left[\partial_{\mu}\left(\mathcal{J}_{\nu}+\epsilon j_{\nu}\right)\right]^{2}+\frac{1}{M_{Z^{\prime}}^{4}}\left(g_{H,2}^{2}+\frac{g^{\prime 2}\epsilon^{2}}{4}\right)\left(H^{\dagger}H\right)\left(\mathcal{J}_{\mu}+\epsilon j_{\mu}\right)^{2}\,,\end{split}$ (12)
@@
$\begin{pmatrix}B_{\mu}\\ Z^{\prime}_{\mu}\end{pmatrix}=\begin{pmatrix}1&\frac{-\epsilon}{\sqrt{1-\epsilon^{2}}}\\ 0&\frac{1}{\sqrt{1-\epsilon^{2}}}\end{pmatrix}\begin{pmatrix}B_{\mu}^{c}\\ Z_{\mu}^{\prime c}\end{pmatrix}\,,$ (13)
@@
$-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}-\frac{\epsilon}{2}B_{\mu\nu}Z^{{}^{\prime}\mu\nu}-\frac{1}{4}Z^{\prime}_{\mu\nu}Z^{{}^{\prime}\mu\nu}\rightarrow-\frac{1}{4}B_{\mu\nu}^{c}B^{c\mu\nu}-\frac{1}{4}Z^{{}^{\prime}c}_{\mu\nu}Z^{{}^{\prime}c\mu\nu}\,.$ (14)
@@ III.1 Dimension-6
$-\frac{1}{8M_{Z^{\prime}}^{2}}\left(2g_{H}+\epsilon g^{\prime}\right)^{2}(H^{\dagger}\overleftrightarrow{D}_{\mu}H)^{2}=-\frac{1}{4M_{Z^{\prime}}^{2}}\left(2g_{H}+\epsilon g^{\prime}\right)^{2}\mathcal{O}_{T},$ (22)
@@
$\frac{C_{\varphi\square}}{\Lambda^{2}}$ $=\frac{1}{8M_{Z^{\prime}}^{2}}\left(2g_{H}+\epsilon g^{\prime}\right)^{2},$ (23)
$\frac{C_{\varphi D}}{\Lambda^{2}}$ $=\frac{1}{2M_{Z^{\prime}}^{2}}\left(2g_{H}+\epsilon g^{\prime}\right)^{2}.$ (24)
Note that since $T\propto C_{T}\propto-\left(2g_{H}+\epsilon g^{\prime}\right)^{2}$, the contribution to the oblique $T$ parameter is always negative. In general, we see that the presence of an $\epsilon$ contribution allows for many directions with vanishing 4-fermion coefficients.
@@ III.2 Dimension-8
Since all operators come from terms of the form $(\mathcal{J}_{\mu}+\epsilon j_{\mu})^{2}$, the operators at dimension-8 may be constructed in terms of the dimension-6 coefficients of the previous section in an obvious way. We define our notation for the generated operators in Table 2, using the notation of Ref. Murphy:2020rsh for operators corresponding to that basis and a superscript $(8)$ for the remainder of the operators. To keep the connection to the dimension-6 operators manifest, we do not attempt to fully translate into the basis of Ref. Murphy:2020rsh . We suppress flavour indices throughout since they are the same between the dimension-8 and dimension-6 pieces. Starting with the four-fermion operators with $H^{\dagger}H$ factors, we find
@@
$\begin{split}\frac{C_{T}^{(8)}}{\Lambda^{4}}&=-\left(\frac{4g_{H,2}^{2}+g^{\prime 2}\epsilon^{2}}{2M_{Z^{\prime}}^{2}}\right)\frac{C_{T}}{\Lambda^{4}},\\ \frac{C_{D^{4}\phi^{4}}^{(8)}}{\Lambda^{4}}&=\left(\frac{1-\epsilon^{2}}{M_{Z^{\prime}}^{2}}\right)\frac{C_{T}}{\Lambda^{2}},\\ \end{split}$ (28)
@@ Acknowledgements.
We would like to thank Yingsheng Huang for helpful correspondence on the results in Ref. Boughezal:2023nhe and Tobias Neumann for assistance with MCFM. We thank Andrea Thamm for an informative seminar on her results on $Z^{\prime}$ bosons. S. D. is supported by the U.S. Department of Energy under Grant Contract DE-SC0012704. M.F. is supported by the U.S. Department of Energy, Office of Science, Office of Workforce Development for Teachers and Scientists, Office of Science Graduate Student Research (SCGSR) program. The SCGSR program is administered by the Oak Ridge Institute for Science and Education (ORISE) for the DOE. ORISE is managed by ORAU under contract number DE-SC0014664. M.S. gratefully acknowledges support from the Alexander von Humboldt Foundation as a Feodor Lynen Fellow. Digital data can be found at https://quark.phy.bnl.gov/Digital_Data_Archive/dawson/zprime_24.

### v3

@@ II Models with a heavy $Z^{\prime}$
$l_{L}^{k}$ $-\frac{1}{2}$ $-1$ $\delta_{ik}-\delta_{jk}$ $\frac{1}{2\sqrt{6}}$ $\frac{3}{2\sqrt{10}}$ $\frac{1}{2\sqrt{15}}$ $-\frac{1}{2}$ $\frac{1}{\sqrt{10}}$ $\frac{2}{\sqrt{15}}$
$\ell^{k}_{R}$ $-1$ $-1$ $\delta_{ik}-\delta_{jk}$ $\frac{1}{2\sqrt{6}}$ $\frac{-1}{2\sqrt{10}}$ $\frac{-1}{\sqrt{15}}$ 00 $\frac{1}{2\sqrt{10}}$ $\frac{-1}{4\sqrt{15}}$
@@ III SMEFT and Matching
$\mathcal{L}_{Z^{\prime}}=-\frac{1}{4}Z^{\prime}_{\mu\nu}Z^{\prime\mu\nu}+\frac{1}{2}M_{Z^{\prime}}^{2}Z^{\prime}_{\mu}Z^{\prime\mu}+\frac{\epsilon}{2}B_{\mu\nu}Z^{\prime\mu\nu}+\left(g_{H,2}\right)^{2}Z^{\prime}_{\mu}Z^{\prime\mu}|H^{\dagger}H|-Z^{\prime}_{\mu}\mathcal{J}^{\mu},$ (4)
@@
where $H^{\dagger}\overleftrightarrow{D}^{\mu}H=H^{\dagger}(D^{\mu}H)-(D^{\mu}H)^{\dagger}H$ with $D_{\mu}=\partial_{\mu}+ig^{\prime}Y_{H}B_{\mu}+igW^{a}_{\mu}T^{a}$ the usual covariant derivative. The couplings of Eqs. 4 and (5) can be determined for each of the models described in the previous section in terms of the $U(1)^{\prime}$ gauge coupling $g_{D}$ and the particle charges of Table 1,
@@
$\begin{split}\delta\mathcal{L}=&-\frac{1}{2M_{Z^{\prime}}^{2}}\mathcal{J}^{\mu}\mathcal{J}_{\mu}+\frac{\epsilon}{M_{Z^{\prime}}^{2}}\left(\partial_{\nu}B^{\mu\nu}\right)\mathcal{J}_{\mu}-\frac{\epsilon^{2}}{2M_{Z^{\prime}}^{2}}\left(\partial_{\nu}B^{\mu\nu}\right)^{2}\\ &+\frac{\epsilon^{2}}{2M_{Z^{\prime}}^{4}}\left(\partial_{\nu}B^{\mu\nu}\right)\partial^{2}\left(\partial^{\alpha}B_{\mu\alpha}\right)-\frac{\epsilon}{M_{Z^{\prime}}^{4}}\left(\partial_{\nu}B^{\mu\nu}\right)\partial^{2}\mathcal{J}_{\mu}+\frac{1}{2M_{Z^{\prime}}^{4}}\mathcal{J}^{\mu}\partial^{2}\mathcal{J}_{\mu}\\ &+\frac{\epsilon^{2}g_{H,2}^{2}}{M_{Z^{\prime}}^{4}}H^{\dagger}H\left(\partial_{\nu}B^{\mu\nu}\right)^{2}-\frac{2\epsilon g_{H,2}^{2}}{M_{Z^{\prime}}^{4}}H^{\dagger}H\left(\partial_{\nu}B^{\mu\nu}\right)\mathcal{J}_{\mu}+\frac{g_{H,2}^{2}}{M_{Z^{\prime}}^{4}}H^{\dagger}H\mathcal{J}^{\mu}\mathcal{J}_{\mu}\,,\end{split}$ (9)
@@
$\begin{split}B_{\mu}&\rightarrow B_{\mu}-\frac{\epsilon^{2}}{2M_{Z^{\prime}}^{2}}\left[\left(\partial^{\nu}B_{\mu\nu}\right)+j_{\mu}\right]+\frac{\epsilon}{M_{Z^{\prime}}^{2}}\mathcal{J}_{\mu}\,,\\ \partial_{\nu}B^{\mu\nu}&\rightarrow\partial_{\nu}B^{\mu\nu}+\frac{\epsilon^{2}}{2M_{Z^{\prime}}^{2}}\partial^{2}\left(\partial_{\nu}B^{\mu\nu}\right)+\frac{\epsilon^{2}}{2M_{Z^{\prime}}^{2}}\partial^{2}j^{\mu}-\frac{\epsilon}{M_{Z^{\prime}}^{2}}\partial^{2}\mathcal{J}^{\mu},\end{split}$ (10)
where $j_{\mu}=-\frac{ig^{\prime}}{2}\left(H^{\dagger}\overleftrightarrow{D^{\mu}}H\right)-g^{\prime}\sum_{f}Y_{f}\bar{f}\gamma^{\mu}f$ is the SM hypercharge current with $f\in(Q_{L},L_{L},u_{R},d_{R},e_{R})$, $g^{\prime}$ is the SM hypercharge gauge coupling and we omit generation indices for simplicity. Note that the definition of $j_{\mu}$ is the same as in Hays:2020scx . Including the newly generated dimension- 8 terms, this gives us
$\begin{split}\delta\mathcal{L}=&-\frac{1}{2M_{Z^{\prime}}^{2}}\mathcal{J}^{\mu}\mathcal{J}_{\mu}-\frac{\epsilon^{2}}{2M_{Z^{\prime}}^{2}}j^{\mu}j_{\mu}+\frac{\epsilon}{M_{Z^{\prime}}^{2}}j^{\mu}\mathcal{J}_{\mu}\\ &+\frac{1}{M_{Z^{\prime}}^{4}}\bigg{\{}\left(\frac{\epsilon^{2}}{2}-\frac{3\epsilon^{4}}{8}\right)\left(\partial_{\nu}B^{\mu\nu}\right)\partial^{2}\left(\partial^{\alpha}B_{\mu\alpha}\right)-\left(\epsilon-\epsilon^{3}\right)\left(\partial_{\nu}B^{\mu\nu}\right)\partial^{2}\mathcal{J}_{\mu}\\ &-\frac{\epsilon^{4}}{4}\left(\partial_{\nu}B^{\mu\nu}\right)\partial^{2}j_{\mu}+\left(\frac{1}{2}-\frac{\epsilon^{2}}{2}\right)\mathcal{J}_{\mu}\partial^{2}\mathcal{J}^{\mu}+\frac{\epsilon^{4}}{8}j^{\mu}\partial^{2}j_{\mu}-\frac{g^{\prime 2}\epsilon^{3}}{4}\left(H^{\dagger}H\right)\mathcal{J}^{\mu}j_{\mu}\\ &+\left(\epsilon^{2}g_{H,2}^{2}+\frac{g^{\prime 2}\epsilon^{4}}{16}\right)\left(H^{\dagger}H\right)\left(\partial_{\nu}B^{\mu\nu}\right)^{2}-\left(2\epsilon g_{H,2}^{2}+\frac{g^{\prime 2}\epsilon^{3}}{4}\right)\left(H^{\dagger}H\right)\left(\partial_{\nu}B^{\mu\nu}\right)\mathcal{J}_{\mu}\\ &+\left(g_{H,2}^{2}+\frac{g^{\prime 2}\epsilon^{2}}{4}\right)\left(H^{\dagger}H\right)\mathcal{J}^{\mu}\mathcal{J}_{\mu}+\frac{g^{\prime 2}\epsilon^{4}}{16}\left(H^{\dagger}H\right)j_{\mu}j^{\mu}+\frac{g^{\prime 2}\epsilon^{4}}{8}\left(H^{\dagger}H\right)\left(\partial_{\nu}B^{\mu\nu}\right)j_{\mu}\bigg{\}}\,.\end{split}$ (11)
@@
$\begin{split}\delta\mathcal{L}=&-\frac{1}{2M_{Z^{\prime}}^{2}}\left(\mathcal{J}_{\mu}-\epsilon j_{\mu}\right)^{2}\\ &-\frac{1}{2M_{Z^{\prime}}^{4}}\left(1-\epsilon^{2}\right)\left[\partial_{\mu}\left(\mathcal{J}_{\nu}-\epsilon j_{\nu}\right)\right]^{2}+\frac{1}{M_{Z^{\prime}}^{4}}\left(g_{H,2}^{2}+\frac{g^{\prime 2}\epsilon^{2}}{4}\right)\left(H^{\dagger}H\right)\left(\mathcal{J}_{\mu}-\epsilon j_{\mu}\right)^{2}\,,\end{split}$ (12)
@@
$\begin{pmatrix}B_{\mu}\\ Z^{\prime}_{\mu}\end{pmatrix}=\begin{pmatrix}1&\frac{\epsilon}{\sqrt{1-\epsilon^{2}}}\\ 0&\frac{1}{\sqrt{1-\epsilon^{2}}}\end{pmatrix}\begin{pmatrix}B_{\mu}^{c}\\ Z_{\mu}^{\prime c}\end{pmatrix}\,,$ (13)
@@
$-\frac{1}{4}B_{\mu\nu}B^{\mu\nu}+\frac{\epsilon}{2}B_{\mu\nu}Z^{{}^{\prime}\mu\nu}-\frac{1}{4}Z^{\prime}_{\mu\nu}Z^{{}^{\prime}\mu\nu}\rightarrow-\frac{1}{4}B_{\mu\nu}^{c}B^{c\mu\nu}-\frac{1}{4}Z^{{}^{\prime}c}_{\mu\nu}Z^{{}^{\prime}c\mu\nu}\,.$ (14)
@@ III.1 Dimension-6
$\frac{1}{8M_{Z^{\prime}}^{2}}\left(2g_{H}+\epsilon g^{\prime}\right)^{2}(H^{\dagger}\overleftrightarrow{D}_{\mu}H)^{2}=\frac{1}{4M_{Z^{\prime}}^{2}}\left(2g_{H}+\epsilon g^{\prime}\right)^{2}\mathcal{O}_{T},$ (22)
@@
$\frac{C_{\varphi\square}}{\Lambda^{2}}$ $=-\frac{1}{8M_{Z^{\prime}}^{2}}\left(2g_{H}+\epsilon g^{\prime}\right)^{2},$ (23)
$\frac{C_{\varphi D}}{\Lambda^{2}}$ $=-\frac{1}{2M_{Z^{\prime}}^{2}}\left(2g_{H}+\epsilon g^{\prime}\right)^{2}.$ (24)
Note that since $T\propto C_{T}\propto\left(2g_{H}+\epsilon g^{\prime}\right)^{2}$, the contribution to the oblique $T$ parameter is always positive. In general, we see that the presence of an $\epsilon$ contribution allows for many directions with vanishing 4-fermion coefficients.
@@ III.2 Dimension-8
Since all operators come from terms of the form $(\mathcal{J}_{\mu}-\epsilon j_{\mu})^{2}$, the operators at dimension-8 may be constructed in terms of the dimension-6 coefficients of the previous section in an obvious way. We define our notation for the generated operators in Table 2, using the notation of Ref. Murphy:2020rsh for operators corresponding to that basis and a superscript $(8)$ for the remainder of the operators. To keep the connection to the dimension-6 operators manifest, we do not attempt to fully translate into the basis of Ref. Murphy:2020rsh . We suppress flavour indices throughout since they are the same between the dimension-8 and dimension-6 pieces. Starting with the four-fermion operators with $H^{\dagger}H$ factors, we find
@@
$\begin{split}\frac{C_{T}^{(8)}}{\Lambda^{4}}&=-\left(\frac{4g_{H,2}^{2}+g^{\prime 2}\epsilon^{2}}{2M_{Z^{\prime}}^{2}}\right)\frac{C_{T}}{\Lambda^{2}},\\ \frac{C_{D^{4}\phi^{4}}^{(8)}}{\Lambda^{4}}&=\left(\frac{1-\epsilon^{2}}{M_{Z^{\prime}}^{2}}\right)\frac{C_{T}}{\Lambda^{2}},\\ \end{split}$ (28)
@@ Acknowledgements.
We would like to thank Yingsheng Huang for helpful correspondence on the results in Ref. Boughezal:2023nhe and Tobias Neumann for assistance with MCFM. We thank Andrea Thamm for an informative seminar on her results on $Z^{\prime}$ bosons. We thank Ben Allenach and Nico Gubernari for pointing out several sign errors in the original version of this paper. S. D. is supported by the U.S. Department of Energy under Grant Contract DE-SC0012704. M.F. is supported by the U.S. Department of Energy, Office of Science, Office of Workforce Development for Teachers and Scientists, Office of Science Graduate Student Research (SCGSR) program. The SCGSR program is administered by the Oak Ridge Institute for Science and Education (ORISE) for the DOE. ORISE is managed by ORAU under contract number DE-SC0014664. M.S. gratefully acknowledges support from the Alexander von Humboldt Foundation as a Feodor Lynen Fellow. Digital data can be found at https://quark.phy.bnl.gov/Digital_Data_Archive/dawson/zprime_24.